1. Evaluate the logarithmic equation for three values of x that are greater than -1, three values of x that are between -2 and -1, and at x = -1.Use the resulting ordered pairs to plot the graph; State the equation of the line asymptotic to the graph (if any).
y = log4 (x + 2)
2. use the formula nkt, where n is the number of items in terms of the initial population i, at time t, and k is the growth constant equal to the percent of growth per unit of time. there are currently 63 million cars in a certain country, decreasing by 6.9% annually. how many years will it take for this country to have 33 million cars? round to the nearest year.

Answer by stanbon(75887)

. Evaluate the logarithmic equation for three values of x that are greater than -1, three values of x that are between -2 and -1, and at x = -1.Use the resulting ordered pairs to plot the graph; State the equation of the line asymptotic to the graph (if any).
y = log4 (x + 2)
---
Comment: Enter the form (log(x+2))/(log(4))into your calculator.
Then find the y values you are looking for under "Table".
-----

2. use the formula nkt, where n is the number of items in terms of the initial population i, at time t, and k is the growth constant equal to the percent of growth per unit of time. there are currently 63 million cars in a certain country, decreasing by 6.9% annually. how many years will it take for this country to have 33 million cars? round to the nearest year.
---
N(t) = No*k^t
33 million = 63 million *(1-0.069)^t
11/21 = 0.931^t
Take the log of both sides to get:
t*log(0.931) = log(11/21)
t = -0.2808/-0.0311
t = 9.0434 years
=======================
Cheers,
Stan H.